Differential Equations Solve: xy + y = 3yPlease name technique used to solve.
Answer #1

zy+ ५.५ = 39 - assume yaar ___(M) + ५. (29) - 2 (9) = 0 (A) - 5,4 - 1 (221)99-2(1-1) 2972-2 (4 - 1) - 391- + ५ = 0 (२- _ ५